# Wenfan Ou and Sheng Bi

### v3

@@ 6 Beyond Rectangularity: Coupled Uncertainty Modelling
While rectangular assumptions bring tractability, such rectangular uncertainty sets are overly conservative in modeling uncertainty and are not always appropriate in practice (Iyengar 2005). The following example demonstrates the benefits of relaxing the rectangular assumption.
@@
A Markov model may be used to describe the health evolution of a patient. The state $s\in\mathcal{S}$ represents the health condition, and the action $a\in A$ represents the treatment. The transition probabilities $\mathbf{p}_{sa}$ represent the dynamics of the health evolution across different health conditions given a treatment. In this context, factors such as genetics and disease traits can influence how patients transition between states, resulting in uncertainty correlation in the transition probabilities across different states.
Inspired by Bertsimas & Sim (2004), Mannor et al. (2012) first propose the coupled ambiguity sets LDST motivated by the famous proverb “lightning does not strike twice”. Concretely, the ambiguity sets consist of support constraints and cardinality constraints which limit the number of times parameters can deviate from the nominal values:
@@
The authors consider two models: a non-adaptive model, which can be viewed as a $T$ single-stage game; and an adaptive model, where both the decision maker and the nature are aware of the number of “deviated visits” over time and modify their strategies. Although the authors provide similar probabilistic guarantees for the two models, solving non-adaptive models is shown to be NP-hard in general and is tractable with only reward uncertainty. In contrast, with the deviation information aware, adaptive models can be tractable via augmenting the state space as
@@
Adhering to the same inspiration, Mannor et al. (2016) further generalizes LDST ambiguity sets to the so-called $k$-rectangular ambiguity sets, treating the LDST ambiguity set as a special case. To ensure universality, the authors introduce the definition of conditional projection which is used for $k$-rectangular ambiguity sets.
@@
Let $\mathcal{S}^{\prime}\subset\mathcal{S}$ be a nonempty subset of the states, the projection of $\mathcal{P}$ to $\mathcal{S}^{\prime}$ is defined as
$\operatorname{Proj}_{\mathcal{S}^{\prime}}\mathcal{P}\triangleq\left\{P_{\mathcal{S}^{\prime}}~{}\left|~{}\exists P_{\mathcal{S}\setminus\mathcal{S}^{\prime}}\colon(P_{\mathcal{S}\setminus\mathcal{S}^{\prime}},P_{\mathcal{S}^{\prime}})\in\mathcal{P}\right\}\right..$ (28)
Given $P_{\mathcal{S}^{\prime}}\in\operatorname{Proj}_{\mathcal{S}^{\prime}}\mathcal{P}$, the projection of $\mathcal{P}$ to $\mathcal{S}\setminus\mathcal{S}^{\prime}$ conditioned on $P_{\mathcal{S}^{\prime}}$ is defined as
$\mathcal{P}_{\mathcal{S}\setminus\mathcal{S}^{\prime}}(P_{\mathcal{S}^{\prime}})\triangleq\{P_{\mathcal{S}\setminus\mathcal{S}^{\prime}}\mid(P_{\mathcal{S}\setminus\mathcal{S}^{\prime}},P_{\mathcal{S}^{\prime}})\in\mathcal{P}\}.$ (29)
Equation (29) could be interpreted as the intersection of $\mathcal{P}$ with affine subspace where $P_{\mathcal{S}^{\prime}}^{0}$ equals $P_{\mathcal{S}^{\prime}}$ from geometric perspective. Without rectangular assumption, $\mathcal{P}_{\mathcal{S}^{*}}(P_{\mathcal{S}^{\prime}})$ obviously depends on $P_{\mathcal{S}^{\prime}}$ and $\mathcal{P}^{\prime}$ for any $\mathcal{S}^{*}\subseteq\mathcal{S}$. Hence, the $k$-rectangular ambiguity sets can be defined rigorously with conditional projection..
@@
For any $\mathcal{S}^{*}\subseteq\mathcal{S}$, a class of conditional projection sets $\mathfrak{P}_{\mathcal{S}^{*}}$ is defined as
$\mathfrak{P}_{\mathcal{S}^{*}}\triangleq\left\{\mathcal{P}_{\mathcal{S}^{*}}(P_{\mathcal{S}^{\prime}})\colon\forall\mathcal{S}^{\prime}\subseteq\mathcal{S}\setminus\mathcal{S}^{*},\forall P_{\mathcal{S}^{\prime}}\in\operatorname{Proj}_{\mathcal{S}^{\prime}}\mathcal{P}\right\},$ (30)
then we call it $k$-rectangular if $|\mathfrak{P}_{\mathcal{S}^{*}}|\leq k$.
The state sets $\mathcal{S}^{\prime}$ can be viewed as the historical observations, and the ambiguity sets conditioned on any $\mathcal{S}^{\prime}\subseteq\mathcal{S}$ contains no more than $k$ sets, where each subset is corresponding to one deviation. Note that (30) is generalized from (26), both the probabilistic guarantee and tractability proof adhere to Mannor et al. (2012) in a similar way.
Different from restricting the cardinality of deviations, Wiesemann et al. (2013) resort to “linear decision rules”, an approach adopted in the RO literature (Ben-Tal et al. 2009), using constant or (piece-wise) linear functions to approximate the value-to-go function. For a non-rectangular ambiguity set $\mathcal{P}$, the authors first construct $\mathcal{\bar{P}}:=\bigotimes_{s\in\mathcal{S}}\mathcal{P}_{s}$ as the smallest $s$-rectangular ambiguity set that contains $\mathcal{P}$. Then, they propose an algorithm to optimize the linear approximations of the value-to-go function over the rectangularized ambiguity set and yield an optimal but overly conservative solution in polynomial time.
Motivated by Example 6.1, Goh et al. (2018) and Goyal & Grand-Clément (2023) emphasize the transition probability is always affected by environment features. Thanks to big data technology, the authors can develop a new framework to explicitly characterize the relationship between features/factors and transition probability, and they call the modeling framework a factor uncertainty model. A factor matrix ambiguity set is defined as
@@
Tirinzoni et al. (2018) provide a distinct scope as a robust improvement. The authors consider a new policy $\pi$, which can achieve robust improvement compared to base policy $\tilde{\pi}$. Being aware that state-action pairs are inefficient in characterizing the underground truth under a non-rectangular setting, they consider incorporating some crucial triple $(s,a,s^{\prime})$ into ambiguity, which can be thought moment information about these important transitions. Concretely, the authors define a feature function $\phi(\cdot):\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$, inspired from inverse reinforcement learning literature, which is always chosen as an indicator function. And then, they construct ambiguity sets as
@@
$\mathcal{P}\triangleq\left\{\mathbf{p}\in(\Delta^{|\mathcal{S}|})^{|\mathcal{S}|\times|\mathcal{A}|}\left|~{}\mathbb{E}_{\mathbf{p}}^{\tilde{\pi}}\left[\textstyle\sum\limits_{t=1}^{T}\phi(s_{t},a_{t},s_{t+1})\right]=\frac{1}{N}\textstyle\sum\limits_{i=1}^{N}\textstyle\sum\limits_{t=1}^{T}\phi(s_{t}^{(i)},a_{t}^{(i)},s_{t+1}^{(i)})\right\}\right..$ (33)
The change from $(s,a)$ pair to $(s,a,s^{\prime})$ triple essentially involves the comprehensive incorporation of entire trajectories.
Recently, Wang, Ho & Petrik (2023) propose a double-loop algorithm for RMDPs, where the outer-loop convergence does not rely on any rectangular assumption. However, in their framework, it is still difficult to solve without any rectangular assumptions for the inner loop.
7 Other Related Literature and New fashion RMDPs
@@
What we have introduced so far still falls within the traditional min-max (max-min) framework, emphasizing the pursuit of an optimal solution to hedge against the (conditional) worst-case. However, the worst-case performance of these approaches is sometimes overly conservative, offering little practical insight into the actual performance of reliable policies. In the RMDPs community, new fashionable modeling approaches have emerged, most inspired by recent advancements in RO and DRO. We will introduce these emerging frameworks and related literature as an extension for future research directions.
@@ 7.1 Satisfactory Regime
RMDPs can mitigate the impact of perturbations, however, the optimized worst-case performances are often too pessimistic and not easy to offer insights and interpretation for decision-makers (Long et al. 2023). To address this issue, the satisfactory RMDPs have emerged. In the satisfactory regime, there are two mainstream formulations: Safe Policy Improvement (SPI) and Robust Satisfying (RS).
SPI is established in Batch Reinforcement Learning (Batch RL), where only limited trajectories and a behavioral/baseline policy $\pi_{B}$ to collect the trajectories are accessible. The goal of SPI is obtaining a policy that is guaranteed to perform at least as well as $\pi_{B}$. The general SPI problem can be described as follows.
@@
Given the uncertainty set $\mathcal{P}$ and a baseline policy $\pi_{B}$, find a maximal $\theta>0$ such that there exists a policy $\pi$ for which $\rho(\pi,\mathbf{p})\geq\rho(\pi_{B},\mathbf{p})+\theta,~{}\forall\mathbf{p}\in\mathcal{P}$, which is equivalent to the following optimization problem:
$\pi_{S}\in\arg\max_{\pi}\min_{\mathbf{p}\in\mathcal{P}}\left[\rho(\pi,\mathbf{p})-\rho(\pi_{B},\mathbf{p})\right].$ (34)
Seminar paper by Ghavamzadeh et al. (2016) considers $L_{1}$-norm uncertainty sets where the discrepancy is bounded by an error function $e(s,a)$. The authors prove that problem (34) may be purely randomized and is an NP-hard problem, even though the uncertainty set is $(s,a)$-rectangular. Due to the intractable formulation, the authors propose a simple and practical approximate algorithm in which either improvement is significant or deviation (error function) is small; otherwise, the baseline policy $\pi_{B}$ is recommended in all other cases.
Subsequently, Laroche et al. (2019) and Simão et al. (2020) coincidentally generalize SPI as SPI with baseline bootstrapping to mitigate the unreliability due to the large state space and restrictive assumptions. The core procedure is to limit state-action pairs into a bootstrapped set $\mathcal{B}$, instead of searching for the analytic optimum over the whole space. In virtue of the statistical properties of bootstrapping, Laroche et al. (2019) proposed a computationally efficient algorithm that provides a PAC-style guarantee of policy improvement with high confidence, and Simão et al. (2020) provide SPI guarantees over the true baseline even without direct access to it.
@@
RS is a novel framework developed in RO and DRO (Long et al. 2023) that transforms the traditional optimization problem of seeking the optimum into attaining an acceptable target specified ahead. Due to the outstanding interpretability and complete theory of RS, Ruan et al. (2023) extend it into the MDPs version in recent. As opposed to traditional RMDPs with hard constraints to limit the distinction among distributions, RS MDPs impose soft constraints for all other transition distributions. Leveraging the theorems from Long et al. (2023), the authors formulate the original problems as a conic program.
The decision variables reflect the magnitude of constraint violation incurred by the distance between the ambiguous transition kernel and the empirical one. Besides, they transform the original problems into a min-max form and propose a primal-dual algorithm to solve the min-max problem efficiently.

### v4

@@ 6 Beyond Rectangularity: Coupled Uncertainty Modeling
While rectangular assumptions bring tractability to RMDPs, such rectangular uncertainty sets are overly conservative in modeling uncertainty and are not always appropriate in practice (Iyengar 2005). The following example demonstrates the necessity of relaxing the rectangular assumptions.
@@
A Markov model may be used to describe the health evolution of a patient. The state $s\in\mathcal{S}$ represents the health condition, and the action $a\in A$ represents the treatment. The transition probabilities $\mathbf{p}_{sa}$ represent the dynamics of the health evolution across different health conditions given a treatment. In this context, factors such as genetics and disease traits can influence how patients transition between states, resulting in uncertainty correlation in the transition probabilities across different states. For example, it is reasonable to believe that two patients in the same moderate state may have different transition probabilities based on their previous conditions: one transitioning from mild to moderate (indicating a worsening trend) and the other from severe to moderate (indicating a recovery trend).
Mannor et al. (2012) first propose the coupled uncertainty sets LDST motivated by the famous proverb “Lightning Does not Strike Twice”. As the proverb goes, the motivation behind LDST is that if the parameters of each state deviate with a small probability, and all states are independent, then the total number of states with deviated parameters will be small. Concretely, the uncertainty sets consist of support constraints and cardinality constraints which limit the number of times parameters can deviate from the nominal values by a given parameter $K$, i.e., at most $K$ states can have transition probabilities different from their nominal estimates:
@@
The authors consider two models: a non-adaptive model, which can be viewed as a single-stage game over $T$ time steps yet with a joint cardinality constraint; and an adaptive model, where both the decision maker and the nature are aware of the number of “deviated visits” over time and modify their strategies dynamically. While similar probabilistic guarantees are provided for both models, the authors demonstrate that solving the non-adaptive model is generally NP-hard, as shown through a reduction from the Vertex Cover Problem. However, the non-adaptive model becomes tractable when limited to reward uncertainty. In contrast, with the deviation information, adaptive models can be tractable by augmenting the state space as
@@
Adhering to the same inspiration, Mannor et al. (2016) further generalizes LDST ambiguity sets to the so-called $k$-rectangular ambiguity sets, treating the LDST ambiguity set as a special case. To ensure readability and completeness, we introduce the definition of conditional projection which is used for $k$-rectangular ambiguity sets.
@@
Let $\mathcal{S}^{\prime}\subset\mathcal{S}$ be a nonempty subset of the states. For a subset $\mathcal{S}^{\prime}\subset\mathcal{S}$, we denote $\mathbf{p}_{\mathcal{S}^{\prime}}=\bigotimes_{s\in\mathcal{S}^{\prime}}\mathbf{p}_{s}$. The projection of an uncertainty set $\mathcal{P}$ to $\mathcal{S}^{\prime}$ is defined as,
$\operatorname{Proj}_{\mathcal{S}^{\prime}}\mathcal{P}\triangleq\left\{\mathbf{p}_{\mathcal{S}^{\prime}}~{}\left|~{}\exists~{}\mathbf{p}_{\mathcal{S}\setminus\mathcal{S}^{\prime}}\colon~{}(\mathbf{p}_{\mathcal{S}\setminus\mathcal{S}^{\prime}},\mathbf{p}_{\mathcal{S}^{\prime}})\in\mathcal{P}\right\}\right..$ (45)
Given $\mathbf{p}_{\mathcal{S}^{\prime}}\in\operatorname{Proj}_{\mathcal{S}^{\prime}}\mathcal{P}$ in (45), the conditional projection of $\mathcal{P}$ to $\mathcal{S}\setminus\mathcal{S}^{\prime}$ with respect to $\mathbf{p}_{\mathcal{S}^{\prime}}$ is defined as
$\mathcal{P}_{\mathcal{S}\setminus\mathcal{S}^{\prime}}(\mathbf{p}_{\mathcal{S}^{\prime}})\triangleq\{\mathbf{p}_{\mathcal{S}\setminus\mathcal{S}^{\prime}}\mid(\mathbf{p}_{\mathcal{S}\setminus\mathcal{S}^{\prime}},\mathbf{p}_{\mathcal{S}^{\prime}})\in\mathcal{P}\}.$ (46)
In essence, the conditional projection $\mathcal{P}_{\mathcal{S}\setminus\mathcal{S}^{\prime}}(\mathbf{p}_{\mathcal{S}^{\prime}})$ in (46) contains all possible parameters (i.e., transition probabilities) of state space $\mathcal{S}\setminus\mathcal{S}^{\prime}$ given $\mathbf{p}_{\mathcal{S}^{\prime}}$, and (45) ensures $\mathbf{p}_{\mathcal{S}^{\prime}}$ is well-defined.
From a geometric perspective, the conditional projection can be viewed as the intersection of $\mathcal{P}$ with an affine subspace where $\mathbf{p}_{\mathcal{S}^{\prime}}$ equals the underlying true parameters $\mathbf{p}^{0}_{\mathcal{S}^{\prime}}$. In other words, the decision-maker can trust the probabilistic information with respect to the state (sub-)space $\mathcal{S}^{\prime}$.
If the uncertainty set $\mathcal{P}$ is rectangular, for any $\mathcal{S}^{*}\subseteq\mathcal{S}$, we readily know $\mathcal{P}_{\mathcal{S}^{*}}$ is unrelated to $\mathbf{p}_{\mathcal{S}^{\prime}}$ when $\mathcal{S}^{*}$ and $\mathcal{S}^{\prime}$ are disjoint, namely, we have $\mathcal{P}=\mathcal{P}_{\mathcal{S}\setminus(\mathcal{S}^{\prime}\cup\mathcal{S}^{*})}\times\mathcal{P}_{\mathcal{S}^{*}}\times\mathcal{P}_{\mathcal{S}^{\prime}}$ by the definition of rectangularity. Conversely, without rectangular assumptions, we cannot obtain such independence. The idea of the $k$-rectangular ambiguity sets is limiting the number of possible conditional projections of $\mathcal{S}^{*}$ given $\mathbf{p}_{\mathcal{S}^{\prime}}$.
@@
For any $\mathcal{S}^{*}\subseteq\mathcal{S}$, let $\mathcal{S}^{\prime}$ range over all subsets of $\mathcal{S}\setminus\mathcal{S}^{*}$ and $\mathbf{p}_{\mathcal{S}^{\prime}}\in\operatorname{Proj}_{\mathcal{S}^{\prime}}\mathcal{P}$, a class of conditional projection sets $\mathfrak{P}_{\mathcal{S}^{*}}$ with respect to $\mathcal{S}^{*}$ is defined as
$\mathfrak{P}_{\mathcal{S}^{*}}\triangleq\left\{\mathcal{P}_{\mathcal{S}^{*}}(\mathbf{p}_{\mathcal{S}^{\prime}})\colon\forall\mathcal{S}^{\prime}\subseteq\mathcal{S}\setminus\mathcal{S}^{*},\forall\mathbf{p}_{\mathcal{S}^{\prime}}\in\operatorname{Proj}_{\mathcal{S}^{\prime}}\mathcal{P}\right\}.$ (47)
The ambiguity set $\mathcal{P}$ is called $k$-rectangular if $|\mathfrak{P}_{\mathcal{S}^{*}}|\leq k$ for all possible subsets $\mathcal{S}^{*}\subseteq\mathcal{S}$.
As we discussed above, we can view the given information $\mathbf{p}_{\mathcal{S}^{\prime}}$ as trustful, and the deviations may occur on the space $\mathcal{S}^{*}$. Limiting the cardinality of the conditional projection sets reduces the possible scenarios. For example, a LDST set constrained by cardinality number $k$ is $(k+1)$-rectangular. Note that (47) is generalized from (43), both the probabilistic guarantee and tractability proof adhere to Mannor et al. (2012) in a similar way.
Different from restricting the cardinality of deviations, Wiesemann et al. (2013) resort to linear decision rules, an approach adopted in the RO literature (Ben et al., 2009), using constant or (piece-wise) linear functions to approximate the value-to-go function. For a non-rectangular ambiguity set $\mathcal{P}$, the authors first construct $\mathcal{\bar{P}}:=\bigotimes_{s\in\mathcal{S}}\mathcal{P}_{s}$ as the smallest $s$-rectangular ambiguity set that contains $\mathcal{P}$. Then, they propose an algorithm to optimize the linear approximations of the value-to-go function over the rectangularized ambiguity set and yield an optimal but overly conservative solution in polynomial time.
Motivated by Example 6.1, Goh et al. (2018) and Goyal & Grand-Clément (2023) argue that the transition probability is always affected by environment features. Thanks to big data technology, the authors can develop a new framework to explicitly characterize the relationship between features/factors and transition probability, and they call this modeling framework a factor uncertainty model. A factor matrix ambiguity set is defined as
@@
As state-action pairs $(s,a)$ are inefficient to characterize the true dynamics in non-rectangular settings, Tirinzoni et al. (2018) propose a robust learning algorithm with policy conditioned marginal uncertainty sets, which incorporates the critical transitions into the construction of uncertainty sets. Specifically, inspired by inverse reinforcement learning (IRL) literature, the authors construct an empirical sample statistics $\hat{\kappa}$ to capture the features of transition tuples $(s,a,s^{\prime})$, instead of directly establishing an empirical distribution as the nominal transition probabilities as same as previous works:
$\hat{\kappa}=\frac{1}{N}\sum\limits_{i=1}^{N}\sum\limits_{t=1}^{T}\phi\left(s_{t}^{(i)},a_{t}^{(i)},s_{t+1}^{(i)}\right)$ (50)
where $\phi(\cdot):\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ is a predetermined feature function. Note that the trajectory samples are generated by a known policy $\tilde{\pi}$, called base policy, under the unknown underlying dynamics $\mathbf{p}^{0}$. The feature function $\phi$ essentially evaluates the “potential value” of each transition $(s,a,s^{\prime})$ induced by $\tilde{\pi}$ and $\mathbf{p}^{0}$. In practice, the feature function $\phi$ is often chosen as an indicator function. In this case, a transition $(s,a,s^{\prime})$ evaluated as 1 represents a crucial event, while those evaluated as 0 can be ignored. This setup can be conceptually linked to LDST or $k$-rectangular set. The empirical statistics $\hat{\kappa}$ approximates the feature expectation $\kappa_{\phi}(\tilde{\pi},\mathbf{p}^{0}):=\mathbb{E}_{\mathbf{p}^{0}}^{\tilde{\pi}}\left[\sum_{t=1}^{T-1}{\phi}(S_{t},A_{t},S_{t+1})\right]$. Given $\hat{\kappa}$, they construct the uncertainty sets as
@@
$\mathcal{P}\triangleq\left\{\mathbf{p}\in(\Delta^{|\mathcal{S}|})^{|\mathcal{S}|\times|\mathcal{A}|}~{}\middle|~{}\kappa_{\phi}(\tilde{\pi},\mathbf{p})=\hat{\kappa}\right\},$ (51)
or a more general form as
$\mathcal{P}\triangleq\left\{\mathbf{p}\in(\Delta^{|\mathcal{S}|})^{|\mathcal{S}|\times|\mathcal{A}|}~{}\middle|~{}\|\kappa_{\phi}(\tilde{\pi},\mathbf{p})-\hat{\kappa}\|\leq\theta\right\}.$ (52)
The uncertainty sets in (51) or (52) require a feasible transition probability matrix $\mathbf{p}$ performs as same as the unknown $\mathbf{p}^{0}$ in terms of feature expectation with feature function $\phi$ and base policy $\tilde{\pi}$. The change from $(s,a)$ pair to $(s,a,s^{\prime})$ triple essentially involves the comprehensive incorporation of entire trajectories. Finally, the authors demonstrate that a robust control problem with a mixed objective with policy conditioned marginal uncertainty sets can be solved in polynomial time. Note that the new issue is not equivalent to the original objective of RMDPs, however, the novel formulation still offers a robust solution in a broad sense.
Recently, Wang et al. (2018) proposed a double-loop algorithm for generic RMDPs, where the outer loop updates the policies and the inner loop updates the worst-case transition probabilities, alternately. In the outer loop, for a policy $\pi_{t}$ at iteration $t$, the algorithm finds a transition $\mathbf{p}_{t}$ of the worst-case one such that
$\mathcal{J}(\pi_{t},\mathbf{p}_{t})\geq\max_{\mathbf{p}\in\mathcal{P}}\mathcal{J}(\pi_{t},\mathbf{p})-\epsilon_{t}$
holds, where $\mathcal{J}(\pi,\mathbf{p}):=\mathbb{E}^{\pi}_{\mathbf{p}}\left[\textstyle\sum\limits_{t=1}^{\infty}\gamma^{t-1}R(s_{t},a_{t},s_{t+1})\right]$ denotes the cumulative discounted reward under the policy $\pi$ and transition matrix $\mathbf{p}$, and $\epsilon_{t}$ is the predetermined tolerance parameter that satisfies $\epsilon_{t+1}\leq\gamma\epsilon_{t}$. Once $\mathbf{p}_{t}$ is determined, it is considered as the worst-case transition probabilities, and a projected gradient step is taken with respect to $\pi$ to minimize $\mathcal{J}(\pi,\mathbf{p}_{t})$ like in non-robust MDPs, thereby obtaining optimized policy $\pi_{t+1}$. A significant contribution is that the outer loop procedure does not require any rectangular assumption, while the procedure for finding $\mathbf{p}_{t}$ is not straightforward. Similarly, the inner loop aims to find the worst-case $\mathbf{p}$ given the fixed outer policy $\pi_{t+1}$. However, the inner maximization remains challenging due to its non-convexity, coinciding with previous results that the policy evaluation procedure is computationally difficult (Wiesemann et al. 2013). To address these challenges, the authors introduce rectangular assumptions specifically for the inner loop in this paper.
7 Other Related Frameworks and New Fashions in RMDPs
@@
What we have introduced so far still falls within the traditional min-max framework, emphasizing the pursuit of an optimal solution to hedge against the (conditional) worst-case. However, the worst-case performance of these approaches is sometimes overly conservative, offering little practical insight into the actual performance of reliable policies. In the RMDPs community, new fashionable modeling approaches have emerged, most inspired by recent advancements in RO and DRO. We will introduce these emerging frameworks and related literature as an extension for future research directions.
@@ 7.1 Satisfactory Regime
RMDPs can mitigate the impact of perturbations, however, the optimized worst-case performances are often too pessimistic and not easy to offer insights and interpretation for decision-makers (Long et al., 2023). To address this issue, the satisfactory RMDPs have emerged. In the satisfactory regime, there are two mainstream formulations: Safe Policy Improvement (SPI) and Robust Satisficing (RS).
SPI is established in Batch Reinforcement Learning, where only limited trajectories and a behavioral/baseline policy $\pi_{B}$ to collect the trajectories are accessible. The goal of SPI is obtaining a policy that is guaranteed to perform at least as well as $\pi_{B}$. The general SPI problem can be described as follows.
@@
Given the uncertainty set $\mathcal{P}$ and a baseline policy $\pi_{B}$, find a maximal $\theta>0$ and a new policy $\pi$ such that $\mathcal{J}(\pi,\mathbf{p})\geq\mathcal{J}(\pi_{B},\mathbf{p})+\theta,~{}\forall\mathbf{p}\in\mathcal{P}$, where $\mathcal{J}(\pi,\mathbf{p})$ denotes the cumulative discounted reward under the policy $\pi$ and transition matrix $\mathbf{p}$. The above statement is mathematically equivalent to the following optimization problem:
$\pi_{S}\in\arg\max_{\pi}\min_{\mathbf{p}\in\mathcal{P}}\left[\mathcal{J}(\pi,\mathbf{p})-\mathcal{J}(\pi_{B},\mathbf{p})\right].$ (53)
Compared to traditional max-min formulation, the problem related to (53) aims to find a policy that maximizes the minimum improvement over the baseline policy $\pi_{B}$, considering all possible realizations in the uncertainty set. It is termed as “safe” because it guarantees an improvement (as $\theta>0$) even in the worst-case scenario within the defined uncertainty set.
The seminal paper in SPI by Ghavamzadeh et al. (2016) considers $L_{1}$-norm uncertainty sets where the discrepancy is bounded by an error function $e(s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$:
$\mathcal{P}_{sa}\triangleq\left.\left\{\mathbf{p}_{sa}\in\Delta^{|\mathcal{S}|}\right|\|\hat{\mathbf{p}}_{sa}-\mathbf{p}_{sa}\|_{1}\leq e(s,a)\right\}.$ (54)
The authors prove that the optimal solution to the problem (53) may be purely randomized by constructing a counterexample, and demonstrate that solving (53) is NP-hard, even though the uncertainty set is $(s,a)$-rectangular. Fortunately, they find that, if the error function induced by the baseline policy is zero, i.e., $e(s,\pi_{B}(s))=0$ for all $s\in\mathcal{S}$, the original problem (53) can be simplified and solved in polynomial time. Motivated by this finding, the authors propose a simple and practical approximate algorithm in which the error function is updated as
$\tilde{e}(s,a)=\begin{cases}e(s,a),&\text{when }\pi_{\mathrm{B}}(s)\neq a\\ 0,&\mathrm{otherwise}\end{cases}\quad\forall~{}(s,a)\in\mathcal{S}\times\mathcal{A}.$
Subsequently, Laroche et al. (2019) and Simao et al. (2020) both generalize SPI as SPI with baseline bootstrapping, effectively mitigating the unreliability caused by the large state space and insufficient samples. The key procedure to achieve this is distinguishing the types of state-action pairs. Specifically, when a state-action pair $(s,a)$ is rarely seen in the dataset (indicating high uncertainty), the trained policy $\pi$ will default to the baseline policy $\pi_{B}$ such that $\pi(a|s)=\pi_{B}(a|s)$. One can see it as a knows-what-it-knows algorithm, seeking assistance from the baseline policy when uncertain, instead of searching for the analytic optimum over the whole space. Particularly, while Laroche et al. (2019) assume the baseline policy $\pi_{B}$ is known, Simão et al. (2020) further relax this assumption and just assume there is an MLE estimate for $\pi_{B}$. Leveraging the statistical properties of bootstrapping, Laroche et al. (2019) propose an efficient algorithm that provides PAC-style guarantees of policy improvement with high probability, and Simão et al. (2020) offer SPI guarantees over the true baseline $\pi_{B}$ even without direct access to it.
@@
Robust Satisficing (RS) is a novel framework developed in RO and DRO (Long et al. 2023) that transforms the traditional optimization problem of seeking the optimum into attaining an acceptable target specified ahead. Due to the outstanding interpretability and solid theoretical foundations of RS, Ruan et al. (2013) extend it into the MDPs version recently.
Recall that solving an MDP can be formulated as an LP. To deliver the core idea of RS, we compare the robust version of dual LP and RS version of dual LP, omitting the corresponding primal form for simplicity. The robust version of dual LP can be computed as
$\begin{array}[]{ll}\max&\bm{r}^{\top}\bm{u}\\ \mathrm{s.t.}&\sum\limits_{a\in\mathcal{A}}u_{sa}-\gamma\sum\limits_{s^{\prime}\in\mathcal{S}}\sum\limits_{a\in\mathcal{A}}p_{s^{\prime}as}u_{s^{\prime}a}-d_{s}\leq 0~{}\forall~{}\mathbf{p}\in\mathcal{P},s\in\mathcal{S}\\ \end{array}$ (55)
where $\bm{r}=(R(s,a))_{s\in\mathcal{S},a\in\mathcal{A}}$ are the immediate rewards and $u_{sa}$ is the dual variable with respect to state-action pair $(s,a)$. Here, uncertainty set $\mathcal{P}=\{\mathbf{p}\in(\Delta^{|\mathcal{S}|})^{|\mathcal{S}|\times|\mathcal{A}|}\mid\ell(\mathbf{p},\hat{\mathbf{p}})\leq\theta\}$. As opposed to traditional RMDPs with hard constraints to limit the distinction among distributions, a target-oriented RSMDP imposes soft constraints for all other transition distributions:
$\begin{array}[]{ll}\min&\sum_{s\in\mathcal{S}}k_{s}\\ \mathrm{s.t.}&\sum\limits_{a\in\mathcal{A}}u_{sa}-\gamma\sum\limits_{s^{\prime}\in\mathcal{S}}\sum\limits_{a\in\mathcal{A}}p_{s^{\prime}as}u_{s^{\prime}a}-d_{s}\leq k_{s}\cdot\ell(\mathbf{p},\hat{\mathbf{p}})~{}\forall~{}s\in\mathcal{S}\\ &\sum\limits_{s\in\mathcal{S}}p_{sa}=1~{}\forall~{}s\in\mathcal{S}\\ &\bm{r}^{\top}\bm{u}\geq\tau\end{array}$ (56)
where $\tau$ is the pre-specified target. To interpret the meanings of the decision variables $k_{s}$ in (56), we introduce the concept of constraint violation, where the constraints in (55) do not hold. Hence, the decision variables $\{k_{s}\}$ reflect the magnitude of constraint violation incurred by the discrepancy between nominal and true transition probabilities. For instance, if the values of $\{k_{s}\}$ are small enough, then $k_{s}\cdot\ell(\mathbf{p},\hat{\mathbf{p}})$ will also tend to be small such that the only mild violation will occur. However, as the existence of $\tau$, the feasible region of $\{k_{s}\}$ is limited, which in turn to restricts the discrepancy $\ell(\mathbf{p},\hat{\mathbf{p}})$.
Compared to (55) which tries to attain the maximal “rewards”, the RSMDP attempts to minimize the magnitude of the constraint violation while ensuring that the “reward” remains at least $\tau$. This approach underscores the concept of satisficing, prioritizing acceptable outcomes over purely optimal rewards. Leveraging the theorems from Long et al. (2023), the authors formulate the original problems as a conic program when the distance function $\ell$ is chosen as $\ell_{1}$-norm. Furthermore, inspired by the FOM framework proposed by Grand-Clément & Kroer (2021), they transform (56) into a min-max form and apply a PDA to solve the min-max problem efficiently.